OpenStudy (anonymous):

Solve the following system of equations. 2x – 3y + z = –1 3x + 2y + 2z = –1 x – y – 3z = –4

OpenStudy (anonymous):

What method you want to use ?

OpenStudy (anonymous):

Solve the last equation for x. Then use it to eliminate x from the first two. This results in two equations iin y and z you can solve

OpenStudy (anonymous):

i just don't know how to do this...it confuses me, i keep getting the wrong answer. Allamiro, elimination would probably be the best!

OpenStudy (anonymous):

Use matrices, yo. They're so much simpler. But if you're really against it... To eliminate, let's look for similar terms. 2x–3y+z=–1 3x+2y+2z =–1 x–y–3z=–4 Oh wait, it'd be awful nice if 2(x–y–3z)=2(–4) could be added on to 2x–3y+z=–1 Do you follow?

OpenStudy (anonymous):

from the last equation: x=-4+y+3z Plug this in for x in the two first. Let me know what you get.

OpenStudy (anonymous):

You take the first equation 2x – 3y + z = –1 and the third x – y – 3z = –4 multiply 1 * 3 and add it to equation 3 7x-10y = -7 take equation 1 and 2 2x–3y+z=–1 3x+2y+2z =–1 multiply 1 * 2 and subtract it x -8y = 0 you take the results

OpenStudy (anonymous):

you will have 7x-10y = -7 x -8y = 0

OpenStudy (anonymous):

I m not sure about the numbers you may want to check my subtraction and addition but you should get the point following the same method find X and Y

OpenStudy (anonymous):

We mult by 7 the second one and subtract -10Y + 8*7 Y =-7

OpenStudy (anonymous):

Thanks for your help guys, i still haven't came up with an answer, so i think ill just skip it.. i was hoping you could solve it for me so i could use it for future reference

OpenStudy (anonymous):

After all this man you need just to use the calculator wooow

OpenStudy (anonymous):

Ill look it over again, thanks for your help & hard work